# Physics meets Topology: Physics-informed topological neural networks for learning rigid body dynamics

### v1

@@ 5.1 Representation of rigid body systems
We introduce a novel representation for environments comprising $K$ interacting rigid bodies. As previously discussed, current methods typically utilize standard graphs $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where vertices $\mathcal{V}$ represent nodes and edges $\mathcal{E}$ denote connections between them. In these approaches, the mesh $M$ of each object, defined by its node positions $\{\bm{x}_{i}\}_{i=1..N}$ and faces $\{f_{j}\}_{j=1..F}$, is integrated into a global graph $\mathcal{G}$. However, this integration leads to the loss of essential structural information, as standard graphs are unable to explicitly define faces and individual objects. Consequently, the detailed geometries and specific interactions of rigid bodies are not accurately captured, thereby limiting the effectiveness of these conventional graph-based methods.
@@
In contrast, our framework leverages a combinatorial complex $\mathcal{X}^{t}$ to model the entire environment at a given time $t$. This complex enables us to explicitly define a hierarchy of higher-order node groups, referred to as cells $X^{(r)}$. Each cell of rank $r$ possesses associated features $\bm{h}^{(r)}$ and can exchange messages with other groups of the same or different ranks. This higher-order representation enables a more precise encoding of both object structure and interactions, ensuring that no critical geometric or topological information is lost during the modeling process.
@@
Our framework defines five distinct types of cells to comprehensively capture the dynamics and interactions of rigid bodies: nodes $X^{(0)}$, edges $X^{(1)}$, mesh triangles $X^{(2)}$, collision contacts $X^{(3)}$, and objects $X^{(4)}$. The cells $X^{(0)}$ and $X^{(1)}$ correspond directly to the nodes $\mathcal{V}$ and edges $\mathcal{E}$ of a conventional graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. To accurately capture the geometric structure of each object, we introduce $X^{(2)}$ cells to represent individual mesh triangles, and $X^{(4)}$ cells to represent entire objects. Furthermore, to model collisions and enable interactions between individual objects, we incorporate collision contacts $X^{(3)}$. These $X^{(3)}$ cells link two mesh triangles $X^{(2)}$ from different objects when they come within a specified collision radius $d_{c}$. This combinatorial complex representation $X^{t}$, alongside the traditional graph representation $\mathcal{G}^{t}$, is illustrated in Fig. 1.
@@
Node features $\bm{h}^{(0),\text{features}}$ for node $i$ belonging to object $k$ consist of two key components: (1) finite-difference velocities derived from the node’s positional history; and (2) center-mass distances between a node and its object’s center of mass. They are formalized in Eq. 2:
$\bm{h}^{(0),\text{features}}_{i}=[\dot{\bm{x}}_{i}^{t},\lVert\dot{\bm{x}}_{i}^{t}\rVert,\dot{\bm{x}}_{i}^{t-1},\lVert\dot{\bm{x}}_{i}^{t-1}\rVert,\bm{d}^{t_{0}}_{i},\lVert\bm{d}^{t_{0}}_{i}\rVert,\bm{d}^{t}_{i},\lVert\bm{d}^{t}_{i}\rVert]$ (2)
$\dot{\bm{x}}^{t}_{i}=\bm{x}^{t}_{i}-\bm{x}^{t-1}_{i}\qquad\bm{d}_{i}^{t}=\bm{x}^{t}_{i}-\bm{x}^{t}_{k}$
@@
Edge features $\bm{h}^{(1),\text{features}}_{s\rightarrow r}$ between sender node $s$ and receiver node $r$ include the node-to-node distances measured at time $t_{0}$ and current time $t$, leading to Eq. 3:
$\bm{h}^{(1),\text{features}}_{s\rightarrow r}=[\bm{x}^{t_{0}}_{s}-\bm{x}^{t_{0}}_{r},\lVert\bm{x}^{t_{0}}_{s}-\bm{x}^{t_{0}}_{r}\rVert,\bm{x}^{t}_{s}-\bm{x}^{t}_{r},\lVert\bm{x}^{t}_{s}-\bm{x}^{t}_{r}\rVert]$ (3)
where $\bm{x}^{t}_{s}$ and $\bm{x}^{t_{0}}_{s}$ are the positions of sender node $s$ at time $t$ and $t_{0}$, and $\bm{x}^{t}_{r}$ and $\bm{x}^{t_{0}}_{r}$ are the positions of receiver node $r$ at time $t$ and $t_{0}$.
@@
Mesh triangle features $\bm{h}^{(2),\text{features}}$ include the normal vector of each mesh triangle $i$:
$\bm{h}_{i}^{(2),\text{features}}=[\bm{n}_{i}^{t},\lVert\bm{n}_{i}^{t}\rVert]$ (4)
where $\bm{n}_{i}^{t}$ is the normal vector vector of mesh triangle $i$ at time $t$.
@@
Collision contact features $\bm{h}^{(3),\text{features}}_{s\rightarrow r}$ between sender triangle $s$ and receiver triangle $r$ are defined by first computing the closest points $\bm{p}_{s}$ and $\bm{p}_{r}$ on each respective triangle. The collision direction is captured by the relative vector $\bm{p_{s}}-\bm{p_{r}}$. Subsequently, a three-axis local coordinate system is established around the closest points on each triangle, utilizing their three vertices $\{s_{j}\}_{j=1,2,3}$ and $\{r_{j}\}_{j=1,2,3}$. This structured approach, formalized in Eq. 5, ensures precise modeling of collision dynamics by accurately representing the spatial relationships between interacting mesh triangles:
$\bm{h}^{(3),\text{features}}_{s\rightarrow r}=[\bm{p}_{s}^{t}-\bm{p}_{r}^{t},\lVert\bm{p}_{s}^{t}-\bm{p}_{r}^{t}\rVert,[\bm{x}_{s_{j}}^{t}-\bm{p}_{s}^{t},\lVert\bm{x}_{s_{j}}^{t}-\bm{p}_{s}^{t}\rVert]_{j=1,2,3},[\bm{x}_{r_{j}}^{t}-\bm{p}_{r}^{t},\lVert\bm{x}_{r_{j}}^{t}-\bm{p}_{r}^{t}\rVert]_{j=1,2,3}]$ (5)
@@
where $\bm{p}_{s}$ is the closest point to triangle $r$ on triangle $s$, $\bm{p}_{r}$ is the closest point to triangle $s$ on triangle $r$, $\{\bm{x}^{t}_{s_{j}}\}_{j=1,2,3}$ are the positions of all three vertices of triangle $s$ and $\{\bm{x}^{t}_{r_{j}}\}_{j=1,2,3}$ are the positions of all three vertices of triangle $r$.
@@
Object features $\bm{h}^{(4),\text{features}}$ for object $k$ include finite-difference velocities computed at the object’s center of mass. Additionally, a one-hot encoded vector $\bm{b}$ distinguishes between static objects and dynamic objects. Finally, some of the object’s physical parameters are also included:
$\bm{h}^{(4),\text{features}}_{k}=[\dot{\bm{x}}_{k}^{t},\lVert\dot{\bm{x}}_{k}^{t}\rVert,\dot{\bm{x}}_{k}^{t-1},\lVert\dot{\bm{x}}_{k}^{t-1}\rVert,\bm{b},m,c_{1},c_{2}]$ (6)
@@
where $\dot{\bm{x}}^{t}_{k}$ and $\dot{\bm{x}}^{t-1}_{k}$ are the finite-difference velocities at time $t$ and $t-1$, $\bm{b}$ is the one-hot encoded object type, $m$ is the object’s mass, $c_{1}$ the friction coefficient and $c_{2}$ the restitution coefficient.
@@
This combinatorial complex representation $\mathcal{X}^{t}$ is inherently translation-invariant, as none of the cell features $\bm{h}^{(r)}$ include the absolute positions of nodes $\{\bm{x}_{i}\}_{i=1..N}$ or objects $\{\bm{x}_{k}\}_{k=1..K}$. This design ensures a translation-equivariant framework, enhancing the model’s generalization capabilities.
@@ 5.2 Physics-informed message passing
Figure 9: Physics-informed message passing. (a) Enhancing mesh face cells $X^{(2)}$ embeddings with nodes, edges, and object cells; (b) Computing the effect of a collision from two mesh faces $s$ and $r$ on each other; (c) Updating triangles $s$ and $r$ after collisions.
@@
Leveraging neighborhood functions, we construct a message-passing neural architecture operating on the combinatorial complexes $\{\mathcal{X}^{t}\}$. Our architecture follows an Encode-Process-Decode graph_network_simulator paradigm. In the encoding step, the cell features $\bm{h}^{(r),\text{features}}$ are passed through separate Multi-Layer Perceptrons (MLPs), one for each cell rank:
$\bm{h}^{(r)}=\text{MLP}^{\text{encoder}}_{r}(\bm{h}^{(r),\text{features}})\vspace{-0.1cm}$ (7)
@@
The processing step is where sequential message-passing occurs. While the topological representation $\mathcal{X}$ and the neighborhood functions $\mathcal{N}(x)$ offer significant flexibility, theoretically supporting learnable messages between all cell ranks, this approach would introduce unnecessary complexity, computational overhead, and negatively impact the learning efficiency of the network. Therefore, it is essential to carefully constrain message-passing pathways, guiding the information flow to only relevant neighboring cells. To define these pathways, we use physics knowledge about collisions and create a sequential message-passing structure, visible in Fig. 3b.
@@
The message-passing used in HOPNet can be divided into three main parts: (A) enhancing mesh faces with intra-object data to prepare for collisions; (B) processing inter-object collisions using mesh faces; (C) updating individual objects and nodes. These three main parts are outlined in Fig. 3b.
@@
A. Enhancing mesh faces with intra-object data. In this first part, the embeddings of $X^{(2)}$ cells are enhanced with messages from nodes $X^{(0)}$ using $\mathcal{N}_{\searrow,2}(x)$, edges $X^{(1)}$ using $\mathcal{N}_{\searrow,1}(x)$, and object $X^{(4)}$ cells using $\mathcal{N}_{\nearrow,2}(x)$. This first main part integrates critical information into the mesh triangles’ embeddings $X^{(2)}$, which are later used to process the inter-object collisions. It is visually illustrated in Fig. 9a, while the messages are formally defined in Eq. 8, and the aggregation in Eq. 9:
$\bm{m}_{j}^{0\rightarrow 2}=\text{MLP}^{\text{processor}}_{m0\rightarrow 2}(\bm{h}_{j}^{(0)})\quad\bm{m}_{j}^{1\rightarrow 2}=\text{MLP}^{\text{processor}}_{m1\rightarrow 2}(\bm{h}_{j}^{(1)})\quad\bm{m}_{j}^{4\rightarrow 2}=\text{MLP}^{\text{processor}}_{m4\rightarrow 2}(\bm{h}_{j}^{(4)})$ (8)
@@
$\bm{h}^{\prime(2)}_{i}=\text{MLP}^{\text{processor}}_{2}([\bm{h}_{i}^{(2)},\sum_{\forall j\in\mathcal{N}_{\searrow,2}(x_{i}^{(2)})}\bm{m}_{j}^{0\rightarrow 2},\sum_{\forall j\in\mathcal{N}_{\searrow,1}(x_{i}^{(2)})}\bm{m}_{j}^{1\rightarrow 2},\sum_{\forall j\in\mathcal{N}_{\nearrow,2}(x_{i}^{(2)})}\bm{m}_{j}^{4\rightarrow 2}])$ (9)
@@
where $\bm{m}^{0\rightarrow 2}$, $\bm{m}^{1\rightarrow 2}$, and $\bm{m}^{4\rightarrow 2}$ are learnable messages – sent from nodes, edges, and object cells to mesh triangles respectively – and $\bm{h}^{\prime(2)}$ are the new mesh triangle embeddings after this message-passing step.

### v3

@@ 5.1 Representation of rigid body systems
We introduce a novel representation for environments comprising $K$ interacting rigid bodies. As previously discussed, current methods typically utilize standard graphs $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where vertices $\mathcal{V}$ represent nodes and edges $\mathcal{E}$ denote connections between them. In these approaches, the mesh $M$ of each object, defined by its node positions $\{\mathbf{x}_{i}\}_{i=1..N}$ and faces $\{f_{j}\}_{j=1..F}$, is integrated into a global graph $\mathcal{G}$. However, this integration leads to the loss of essential structural information, as standard graphs are unable to explicitly define faces and individual objects. Consequently, the detailed geometries and specific interactions of rigid bodies are not accurately captured, thereby limiting the effectiveness of these conventional graph-based methods.
@@
In contrast, our framework leverages a combinatorial complex $\mathcal{X}^{t}$ to model the entire environment at a given time $t$. This complex enables us to explicitly define a hierarchy of higher-order node groups, referred to as cells $X^{(r)}$. Each cell of rank $r$ possesses associated features $\mathbf{h}^{(r)}$ and can exchange messages with other groups of the same or different ranks. This higher-order representation enables a more precise encoding of both object structure and interactions, ensuring that no critical geometric or topological information is lost during the modeling process.
@@
Our framework defines five distinct types of cells to comprehensively capture the dynamics and interactions of rigid bodies: nodes $X^{(0)}$, edges $X^{(1)}$, mesh triangles $X^{(2)}$, collision contacts $X^{(3)}$, and objects $X^{(4)}$. The cells $X^{(0)}$ and $X^{(1)}$ correspond directly to the nodes $\mathcal{V}$ and edges $\mathcal{E}$ of a conventional graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. To accurately capture the geometric structure of each object, we introduce $X^{(2)}$ cells to represent individual mesh triangles, and $X^{(4)}$ cells to represent entire objects. Furthermore, to model collisions and enable interactions between individual objects, we incorporate collision contacts $X^{(3)}$. These $X^{(3)}$ cells link two mesh triangles $X^{(2)}$ from different objects when they come within a specified collision radius $d_{c}$. This combinatorial complex representation $X^{t}$, alongside the traditional graph representation $\mathcal{G}^{t}$, is illustrated in Figure 1.
5.1.1 Node features
@@
Node features $\mathbf{h}^{(0),\text{features}}$ for node $i$ belonging to object $k$ consist of two key components: (1) finite-difference velocities derived from the node’s positional history; and (2) center-mass distances between a node and its object’s center of mass. They are formalized in Equation 2:
$\mathbf{h}^{(0),\text{features}}_{i}=[\dot{\mathbf{x}}_{i}^{t},\lVert\dot{\mathbf{x}}_{i}^{t}\rVert,\dot{\mathbf{x}}_{i}^{t-1},\lVert\dot{\mathbf{x}}_{i}^{t-1}\rVert,\mathbf{d}^{t_{0}}_{i},\lVert\mathbf{d}^{t_{0}}_{i}\rVert,\mathbf{d}^{t}_{i},\lVert\mathbf{d}^{t}_{i}\rVert]$ (2)
$\dot{\mathbf{x}}^{t}_{i}=\mathbf{x}^{t}_{i}-\mathbf{x}^{t-1}_{i}\qquad\mathbf{d}_{i}^{t}=\mathbf{x}^{t}_{i}-\mathbf{x}^{t}_{k}$
@@ 5.1.2 Edge features
Edge features $\mathbf{h}^{(1),\text{features}}_{s\rightarrow r}$ between sender node $s$ and receiver node $r$ include the node-to-node distances measured at time $t_{0}$ and current time $t$, leading to Equation 3:
$\mathbf{h}^{(1),\text{features}}_{s\rightarrow r}=[\mathbf{x}^{t_{0}}_{s}-\mathbf{x}^{t_{0}}_{r},\lVert\mathbf{x}^{t_{0}}_{s}-\mathbf{x}^{t_{0}}_{r}\rVert,\mathbf{x}^{t}_{s}-\mathbf{x}^{t}_{r},\lVert\mathbf{x}^{t}_{s}-\mathbf{x}^{t}_{r}\rVert]$ (3)
where $\mathbf{x}^{t}_{s}$ and $\mathbf{x}^{t_{0}}_{s}$ are the positions of sender node $s$ at time $t$ and $t_{0}$, and $\mathbf{x}^{t}_{r}$ and $\mathbf{x}^{t_{0}}_{r}$ are the positions of receiver node $r$ at time $t$ and $t_{0}$.
5.1.3 Mesh triangle features
@@
Mesh triangle features $\mathbf{h}^{(2),\text{features}}$ include the normal vector of each mesh triangle $j$:
$\mathbf{h}_{j}^{(2),\text{features}}=[\mathbf{n}_{j}^{t},\lVert\mathbf{n}_{j}^{t}\rVert]$ (4)
where $\mathbf{n}_{j}^{t}$ is the normal vector vector of mesh triangle $j$ at time $t$.
5.1.4 Collision contact features
@@
Collision contact features $\mathbf{h}^{(3),\text{features}}_{s\rightarrow r}$ between sender triangle $s$ and receiver triangle $r$ are defined by first computing the closest points $\mathbf{p}_{s}$ and $\mathbf{p}_{r}$ on each respective triangle. The collision direction is captured by the relative vector $\mathbf{p_{s}}-\mathbf{p_{r}}$. Subsequently, a three-axis local coordinate system is established around the closest points on each triangle, utilizing their three vertices $\{s_{i}\}_{{i}=1,2,3}$ and $\{r_{i}\}_{{i}=1,2,3}$. This structured approach, formalized in Equation 5, ensures precise modeling of collision dynamics by accurately representing the spatial relationships between interacting mesh triangles:
$\mathbf{h}^{(3),\text{features}}_{s\rightarrow r}=[\mathbf{p}_{s}^{t}-\mathbf{p}_{r}^{t},\lVert\mathbf{p}_{s}^{t}-\mathbf{p}_{r}^{t}\rVert,[\mathbf{x}_{s_{i}}^{t}-\mathbf{p}_{s}^{t},\lVert\mathbf{x}_{s_{i}}^{t}-\mathbf{p}_{s}^{t}\rVert]_{{i}=1,2,3},[\mathbf{x}_{r_{i}}^{t}-\mathbf{p}_{r}^{t},\lVert\mathbf{x}_{r_{i}}^{t}-\mathbf{p}_{r}^{t}\rVert]_{{i}=1,2,3}]$ (5)
@@
where $\mathbf{p}_{s}$ is the closest point to triangle $r$ on triangle $s$, $\mathbf{p}_{r}$ is the closest point to triangle $s$ on triangle $r$, $\{\mathbf{x}^{t}_{s_{i}}\}_{{i}=1,2,3}$ are the positions of all three vertices of triangle $s$ and $\{\mathbf{x}^{t}_{r_{i}}\}_{{i}=1,2,3}$ are the positions of all three vertices of triangle $r$.
5.1.5 Object features
@@
Object features $\mathbf{h}^{(4),\text{features}}$ for object $k$ include finite-difference velocities computed at the object’s center of mass. Additionally, a one-hot encoded vector $\mathbf{b}$ distinguishes between static objects and dynamic objects. Finally, some of the object’s physical parameters are also included:
$\mathbf{h}^{(4),\text{features}}_{k}=[\dot{\mathbf{x}}_{k}^{t},\lVert\dot{\mathbf{x}}_{k}^{t}\rVert,\dot{\mathbf{x}}_{k}^{t-1},\lVert\dot{\mathbf{x}}_{k}^{t-1}\rVert,\mathbf{b},m,c_{1},c_{2}]$ (6)
@@
where $\dot{\mathbf{x}}^{t}_{k}$ and $\dot{\mathbf{x}}^{t-1}_{k}$ are the finite-difference velocities at time $t$ and $t-1$, $\mathbf{b}$ is the one-hot encoded object type, $m$ is the object’s mass, $c_{1}$ the friction coefficient and $c_{2}$ the restitution coefficient.
@@
This combinatorial complex representation $\mathcal{X}^{t}$ is inherently translation-invariant, as none of the cell features $\mathbf{h}^{(r)}$ include the absolute positions of nodes $\{\mathbf{x}_{i}\}_{i=1..N}$ or objects $\{\mathbf{x}_{k}\}_{k=1..K}$. This design ensures a translation-equivariant framework, enhancing the model’s generalization capabilities.
HOPNet is a fully autoregressive model: it relies solely on past information from timesteps $t-1$ and $t$ to predict the next state at $t+1$. None of the cell features $\mathbf{h}^{(r),\text{features}}$ in HOPNet incorporate any future information.
@@ 5.2 Physics-informed message passing
Figure 10: Physics-informed message passing. (a) Enhancing mesh face cells $X^{(2)}$ embeddings with nodes, edges, and object cells; (b) Computing the effect of a collision from two mesh faces $s$ and $r$ on each other; (c) Updating triangles $s$ and $r$ after collisions. Here, $\mathbf{h}^{(r)}$ denotes the embedding of a cell of rank $r$, and $\mathbf{m}^{r_{a}\rightarrow r_{b}}$ represents a message passed from a cell of rank $r_{a}$ to one of rank $r_{b}$.
@@
Leveraging neighborhood functions, we construct a message-passing neural architecture operating on the combinatorial complexes $\{\mathcal{X}^{t}\}$. Our architecture follows an Encode-Process-Decode graph_network_simulator paradigm. In the encoding step, the cell features $\mathbf{h}^{(r),\text{features}}$ are passed through separate MLPs, one for each cell rank:
$\mathbf{h}^{(r)}=\text{MLP}^{\text{encoder}}_{r}(\mathbf{h}^{(r),\text{features}})\vskip-2.84544pt$ (7)
@@
The processing step is where sequential message-passing occurs. While the topological representation $\mathcal{X}$ and the neighborhood functions $\mathcal{N}(x)$ offer significant flexibility, theoretically supporting learnable messages between all cell ranks, this approach would introduce unnecessary complexity, computational overhead, and negatively impact the learning efficiency of the network. Therefore, it is essential to carefully constrain message-passing pathways, guiding the information flow to only relevant neighboring cells. To define these pathways, we use physics knowledge about collisions and create a sequential message-passing structure, visible in Figure 3b.
@@
The message-passing used in HOPNet can be divided into three main parts: (A) enhancing mesh faces with intra-object data to prepare for collisions; (B) processing inter-object collisions using mesh faces; (C) updating individual objects and nodes. These three main parts are outlined in Figure 3b.
5.2.1 Enhancing mesh faces with intra-object data
@@
In this first part, the embeddings of $X^{(2)}$ cells are enhanced with messages from nodes $X^{(0)}$ using $\mathcal{N}_{\searrow,2}(x)$, edges $X^{(1)}$ using $\mathcal{N}_{\searrow,1}(x)$, and object $X^{(4)}$ cells using $\mathcal{N}_{\nearrow,2}(x)$. This first main part integrates critical information into the mesh triangles’ embeddings $X^{(2)}$, which are later used to process the inter-object collisions. It is illustrated in Figure 9a, while the messages are formally defined in Equation 8, and the aggregation in Equation 9:
$\mathbf{m}_{j}^{0\rightarrow 2}=\text{MLP}^{\text{processor}}_{m0\rightarrow 2}(\mathbf{h}_{j}^{(0)})\quad\mathbf{m}_{j}^{1\rightarrow 2}=\text{MLP}^{\text{processor}}_{m1\rightarrow 2}(\mathbf{h}_{j}^{(1)})\quad\mathbf{m}_{j}^{4\rightarrow 2}=\text{MLP}^{\text{processor}}_{m4\rightarrow 2}(\mathbf{h}_{j}^{(4)})$ (8)
@@
$\mathbf{h}^{\prime(2)}_{i}=\text{MLP}^{\text{processor}}_{2}([\mathbf{h}_{i}^{(2)},\sum_{\forall j\in\mathcal{N}_{\searrow,2}(x_{i}^{(2)})}\mathbf{m}_{j}^{0\rightarrow 2},\sum_{\forall j\in\mathcal{N}_{\searrow,1}(x_{i}^{(2)})}\mathbf{m}_{j}^{1\rightarrow 2},\sum_{\forall j\in\mathcal{N}_{\nearrow,2}(x_{i}^{(2)})}\mathbf{m}_{j}^{4\rightarrow 2}])$ (9)
@@
where $\mathbf{m}^{0\rightarrow 2}$, $\mathbf{m}^{1\rightarrow 2}$, and $\mathbf{m}^{4\rightarrow 2}$ are learnable messages—sent from nodes, edges, and object cells to mesh triangles respectively—and $\mathbf{h}^{\prime(2)}$ are the new mesh triangle embeddings after this message-passing step.
5.2.2 Processing inter-object collisions using mesh faces